OpenStudy (anonymous):

what is the equation for m: 9m = 108

Nnesha (nnesha):

9 is multiplying with m do cancel out 9 from left side do opposite of "multiply" both side

OpenStudy (anonymous):

ok thx

OpenStudy (kitkat16):

9*m=108 what is m?

OpenStudy (anonymous):

i think it is m=12

OpenStudy (kitkat16):

yes

Nnesha (nnesha):

so opposite of multiplication is division so divide both side by 9 you can check your answer by plugging 12 into the equation

Nnesha (nnesha):

\[\huge\rm 9 \times \color{red}{m }= 108\]\[\color{red}{12} \times 9 =108\] both sides should be equal :-)